OpenStudy (xo.msml):

A pair of equations is shown below: y = 3x − 5 y = 6x − 8 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer. (5 points)

OpenStudy (xo.msml):

@rebeccaxhawaii @sloppycanada

OpenStudy (photon336):

well y = 3x-5 right we can set the first equation equal to the second one

OpenStudy (photon336):

3x-5 = 6x-8 then solve for x

OpenStudy (photon336):

@xo.msml

OpenStudy (xo.msml):

x=1

OpenStudy (photon336):

so plug that back into the first equation to find out what y is

OpenStudy (xo.msml):

-5

OpenStudy (xo.msml):

@Photon336

OpenStudy (photon336):

3(1)-5 = 3-5 = -2 6(1)-8 = 6-8 = ?

OpenStudy (xo.msml):

-2?

OpenStudy (photon336):

yes

OpenStudy (xo.msml):

what do i do next? I am really bad at this

OpenStudy (photon336):

that's actually the solution (1,-2)

OpenStudy (photon336):

that's the point where the graphs intersect.

OpenStudy (xo.msml):

oh wow thank you so much!!! Really appreciate it

OpenStudy (xo.msml):

will it be substitution or elimination

OpenStudy (photon336):

No problem

OpenStudy (photon336):

that's substitution